Calculate the rate at which a black body at temperature 300 degrees Celsius emits radiant energy. The black body is a sphere of diameter 20cm
1
Expert's answer
2020-08-10T12:32:31-0400

According to the Stefan-Bolzmann law,

"L= 4\pi R^2 \sigma T^4 = 4\pi \cdot (0.1)^2 \cdot 5.67\cdot10^{-8}\cdot (273+300)^4 \approx 768\,\mathrm{W}."